P and Q are partners in a venture,P contributed $20,000 for nine months and Q contributed $50,000 for one year:find each person's share of profit of $6,300?
1
Expert's answer
2019-04-08T09:01:33-0400

If P contributed $20,000 for nine months and Q contributed $50,000 for one year, then P contributed 20,000*9 = $180,000 in total and Q contributed 50,000*12 = $600,000 in total. Their total contribution is $780,000, so P contributed 180,000/780,000 = 0.231 or 23.1% and Q contributed 600,000/780,000 = 0.769 or 76.9%.

Each person's share of profit of $6,300 is:

P will receive 0.231*6,300 = $1,453.85,

Q will receive 6,300 - 1,453.85 = $4, 846.15.